# Tensor Product Attention Is All You Need

### v3

@@ 3 Tensor Product Attention
In this section, we provide a detailed description of our proposed Tensor Product Attention (TPA), which allows contextual low-rank factorization for queries, keys, and values. First, we explain how TPA factorizes queries, keys, and values with explicit tensor shapes. Next, we describe how TPA can be integrated into the multi-head attention framework and how it reduces memory consumption in KV caching at inference time. Finally, we show how RoPE can seamlessly integrate with TPA (including a pre-rotated variant).
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Contextual Factorization (CF). Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ. Specifically, for each token $t$, with a small abuse of notation, we define:
$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),$ (3.1)
$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$ (3.2)
$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.3)
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Similarly, analogous definitions apply to key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
and similarly for keys and values.
@@
which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times d_{h}}$. Summing over $R_{q}$ and scaled by $\frac{1}{R_{q}}$ yields
@@
Repeating for all tokens reconstitutes $\mathbf{Q}\in\mathbb{R}^{T\times h\times d_{h}}$. Similar procedures can be applied to obtain $\mathbf{K}$ and $\mathbf{V}$ with ranks $R_{k}$ and $R_{v}$, respectively.
@@
Parameter Initialization. We initialize the weight matrices $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, $\bm{W}_{r}^{b^{V}}$ using Xavier initialization (Glorot & Bengio, 2010). Specifically, each entry of the weight matrix is drawn from a uniform distribution with bounds $[-\sqrt{6/(n_{\text{in}}+n_{\text{out}})},\sqrt{6/(n_{\text{in}}+n_{\text{out}})}]$, where $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrices. This initialization strategy helps maintain the variance of activations and gradients across the network.
@@ 3.2 RoPE Compatibility and Acceleration
$\mathbf{Q}_{t}\mapsto\widetilde{\mathbf{Q}_{t}}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t}),\qquad\mathbf{K}_{s}\mapsto\widetilde{\mathbf{K}_{s}}=\operatorname{RoPE}_{s}(\mathbf{K}_{s}).$
@@
$\widetilde{\mathbf{K}}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}$
$=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
Thus, each $\mathbf{K}_{t}$ is already rotated before caching, removing the need for explicit rotation at the decoding time and accelerating autoregressive inference. Depending on hardware and performance requirements, one can also adopt different RoPE integration approaches for training and inference.
Theorem 1 (RoPE’s Compatibility with TPA).
@@
where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$. In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$. Then we have
@@
Focusing on individual heads $i$, the above matrix equality implies:
$\operatorname{RoPE}_{t-s}\bigl{(}\mathbf{q}_{t,i}\bigr{)}^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}.$
where $\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$ is the $i$-th query head of $t$-th token, and $\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$ is the $j$-th key head of $s$-th token, and
$\widetilde{\mathbf{q}}_{t,i}$ $=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$
$\quad\widetilde{\mathbf{k}}_{s,i}$ $=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
Theorem 1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 1 in Appendix C.1. In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$. Consequently, $\mathbf{A}_{Q}(\mathbf{x}_{t})$ remains unchanged, while each column of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ is rotated appropriately, preserving the TPA structure.
@@ 3.3 KV Caching and Memory Reduction
TPA Factorized KV Caching. Instead of storing the full $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$, TPA stores only their factorized ranks. Specifically, we keep
@@
where $\mathbf{A}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times h},\;\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times d_{h}},\;\mathbf{A}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times h},\;\mathbf{B}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times d_{h}}$.
Hence, the memory cost per token is
$\underbrace{R_{K}(h+d_{h})}_{\text{for K}}\;+\;\underbrace{R_{V}(h+d_{h})}_{\text{for V}}=(\,R_{K}+R_{V}\,)\,\bigl{(}h+d_{h}\bigr{)}.$
Compared to the standard caching cost of $2\,h\,d_{h}$, the ratio is:
$\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}.$
For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section F. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
3.4 Unifying MHA, MQA, and GQA as Non-contextual TPA
3.4.1 MHA as Non-contextual TPA
Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$.
To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$. Focusing on $\mathbf{Q}_{t}$:
  1. (a)
    Non-contextual head factors. Define
    $\mathbf{a}^{Q}_{i}=R_{Q}\mathbf{e}_{i}\in\mathbb{R}^{h},$ (3.8)
  2. (b)
    Contextual token factors. Define
    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (3.9)
    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection defined before, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ dependent on $\mathbf{x}_{t}$.
Substituting (3.8)–(3.9) into (3.1) gives:
$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (3.10)
Each term $\mathbf{e}_{i}\otimes\bigl{(}({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\bigr{)}$ in (3.10) contributes only to the $i$-th row, reconstituting the usual MHA form of $\mathbf{Q}_{t}$. Analogous constructions hold for $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ using $\bm{W}^{K}_{i},\bm{W}^{V}_{i}$. Thus, MHA is a non-contextual, full-rank variant of TPA.
TPA with Non-contextual A. More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent. Then, for keys:
and similarly for queries/values. This reduces per-token computations and can be effective when head-dimension relationships are relatively stable across all tokens.
MQA and GQA as Non-Contextual TPA. Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) [^6] The original definitions of MQA and GQA are presented in Appendix E.1 and E.2, respectively. also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
  • MQA as Rank-1 TPA. In MQA, all heads share a single set of keys/values, corresponding to $R_{K}=R_{V}=1$ along the head dimension. Concretely,
    $\mathbf{K}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),$
    $\mathbf{V}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
    forces every head to use the same $\mathbf{K}_{t},\mathbf{V}_{t}$. Each head retains a distinct query projection, matching the MQA design.
  • GQA as Grouped Rank-1 TPA. GQA partitions $h$ heads into $G$ groups, each sharing keys/values within that group. In TPA form, each group $g$ has a dedicated non-contextual factor pair $\mathbf{a}^{K}_{g},\mathbf{a}^{V}_{g}\in\mathbb{R}^{h}$, which acts as a “mask” for the heads in that group. Varying $G$ from $1$ to $h$ interpolates from MQA to standard MHA.
Hence, by constraining TPA’s head-dimension factors to be constant masks (one for MQA; multiple for GQA), these popular variants are recovered as special cases.
3.5 Computational Cost
For a detailed analysis of the computational cost of TPA, please refer to Appendix A, which shows that the training and inference flops of TPA with optimized implementation (without materializing $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) are smaller than MHA, GQA, and MLA. Specifically, when we set $R_{q}=6$, $R_{k}=R_{v}=2$ (our default setting), TPA is 10×10$\times$10 × or more faster on calculating $\mathbf{Q}\mathbf{K}^{\top}$ than MLA during inference (see Appendix A.8).

### v4

@@ 3 Tensor Product Attention
In this section, we provide a detailed description of our proposed Tensor Product Attention (TPA), which enables contextual low-rank factorization for queries, keys, and values. First, we explain how TPA factorizes these components, specifying tensor shapes. Next, we describe TPA’s integration into the multi-head attention framework and its benefits for reducing KV cache memory consumption during inference. Finally, we demonstrate RoPE’s seamless integration with TPA, including a pre-rotated variant for efficiency.
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Contextual Factorization. Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ. Specifically, for each token $t$, with a small abuse of notation, we define:
[…]
$\hskip 77.49976pt\mathbf{V}_{t}=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.1)
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
where $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$ are learnable weight matrices. Similar linear maps produce the factors for keys and values.
@@
which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times d_{h}}$ (where each row of $\mathbf{A}_{Q}(\mathbf{x}_{t})$ corresponds to an $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})^{\top}$ and each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to a $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})^{\top}$). The query tensor for token $t$ can then be expressed as:
@@
This operation is equivalent to $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})(\mathbf{b}_{r}^{Q}(\mathbf{x}_{t}))^{\top}$, where $\mathbf{a}_{r}^{Q}$ is the $r$-th column of $\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}$ and $(\mathbf{b}_{r}^{Q})^{\top}$ is the $r$-th row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$. Repeating for all tokens reconstitutes $\mathbf{Q}\in\mathbb{R}^{T\times h\times d_{h}}$. Similar procedures are applied to obtain $\mathbf{K}$ and $\mathbf{V}$ with ranks $R_{K}$ and $R_{V}$, respectively.
@@
Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in Appendix I.
@@ 3.2 RoPE Compatibility and Acceleration
$\mathbf{Q}_{t}\mapsto\widetilde{\mathbf{Q}_{t}}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t}),\qquad\mathbf{K}_{s}\mapsto\widetilde{\mathbf{K}_{s}}=\operatorname{RoPE}_{s}(\mathbf{K}_{s}).$ (3.4)
@@
$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})\bigr{)}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})$ vector). Thus, each $\mathbf{K}_{t}$ is effectively rotated before caching. This removes the need for explicit rotation at decoding time, accelerating autoregressive inference. Depending on hardware and performance requirements, different RoPE integration strategies can be adopted for training and inference.
Theorem 3.1 (RoPE’s Compatibility with TPA).
@@
where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$). Furthermore, let $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ be factorized by TPA. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$ be their RoPE-transformed versions. The relative positional encoding property of RoPE is preserved:
@@
where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position $t-s$. Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
$\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}=\mathbf{q}_{t,i}^{\top}\operatorname{RoPE}(\cdot,t-s)\mathbf{k}_{s,i}.$
Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 3.1 in Appendix F.1. In essence, $\operatorname{RoPE}_{t}$ applies a linear transformation (rotation matrix $\mathbf{R}_{t}$) to the $d_{h}$-dimensional space. Since $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, applying RoPE yields $\mathbf{Q}_{t}\mathbf{R}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t})$. Thus, RoPE effectively transforms $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t}$, where $\mathbf{R}_{t}$ acts on each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$. The $\mathbf{A}_{Q}(\mathbf{x}_{t})$ factor remains unchanged, preserving the TPA structure.
@@ 3.3 KV Caching and Memory Reduction
TPA Factorized KV Caching. Instead of storing the full $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$, TPA stores only their factor components. Specifically, for each past token $t$, we cache:
@@
where $\mathbf{A}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times h},\;\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times d_{h}}(\text{pre-rotated}),\;\mathbf{A}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times h},\;\mathbf{B}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times d_{h}}$.
Hence, the memory cost per token is $\underbrace{R_{K}(h+d_{h})}_{\text{for K}}\;+\;\underbrace{R_{V}(h+d_{h})}_{\text{for V}}=(\,R_{K}+R_{V}\,)\,\bigl{(}h+d_{h}\bigr{)}$. Compared to the standard caching cost of $2\,h\,d_{h}$, the ratio is $\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}$. For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields substantial reduction of KV cache size. Table 1 provides a comparative overview of different attention mechanisms, including TPA and its variants, focusing on KV cache size per token and the number of parameters in an attention layer.
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Appendix I. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
3.4 Expressing MHA, MQA, and GQA as Non-contextual TPA
We demonstrate in Appendix A that several existing attention mechanisms, including MHA, MQA, and GQA, can be elegantly expressed as non-contextual variants of TPA by imposing specific constraints on the TPA factors.